If a typical aluminum beverage can weighs 13.5g, how much alum could you produce for every 210 cans collected? Show your work based on your own experiment.

My experiment is Aluminum recycling and here are my results: mass of aluminum can fragments = 0.4559g ; mass of alum produced = 6.2400g ; theoretical yield of alum, 2KA1(SO4)2•12H2O(s) = 8.01498g


1
Expert's answer
2021-09-29T05:01:06-0400

According to your experiment, every 0.4787 g of Aluminium can produce 7.0737 g of alum.

Therefore the mass of alum produced from 1 g of Aluminium is:

"mass\: of\:alum\: produced\: from\: 1g\: Al=\frac{7.0737}{0.4787}=14.7769\: g"

Mass of 1 aluminum can = 13.5 g

Mass of 210 aluminum cans = 210×13.5 g = 2835 g

Mass of alum produced from 1 g of aluminum is 14.7769 g

Mass of alum that can be produced from 2835 g of aluminium = 2835×14.7769 g = 41892.5 g = 41.89 kg

Hence, for every 210 cans collected, you can produce 41.89 kg of alum.